OpenStudy (anonymous):

Two runners start 30 kilometers apart and run toward each other. They meet after 3 hours. One runner runs 2 kilometers per hour faster than the other runner. What rate do they run?

OpenStudy (anonymous):

@Cpt.Sparz

OpenStudy (anonymous):

your asked to find rate, so that's going to be your unknown: let x be the first, slower speed, and x+2 be the second faster speed. each of these rates run for 3 hrs should total 30 kms, so we have \[\large D_{Total}=D_{slow}+D_{fast}\] \[\large 30kms=[Rate_{slow}*time]+[Rate_{fast}*time]\] \[\large 30[kms] = x*3[hrs]+(x+2)*3[hrs]\] solve for x and the units should give you km/hr

OpenStudy (anonymous):

Mathmuse is correct...

OpenStudy (anonymous):

For these questions, distances can be added directly. Rates of speed cannot be added correctly so you usually have generate a formula with distance and break it into components as per the distance/speed/time relationship: 

OpenStudy (anonymous):

But if you want the answer here: The sum of their speeds is (30 km)/(3 hr) = 10 kph. The difference of their speeds is 2 kph (from the problem statement). This has solutions like any other "sum and difference" problem. The faster runner is running (10 + 2)/2 = 6 kph. The slower runner is running (10 - 2)/2 = 4 kph.

OpenStudy (anonymous):

I THINK THIS IS MY ANWERS 4 km/hr and 6 km/hr

OpenStudy (anonymous):

instead of km use kph so 4 kph/hr and 6 kph/hr

OpenStudy (anonymous):

kph=kilometers per hour

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

IS CORRECT MY ANSWER

OpenStudy (anonymous):

yes your answer is right!